# On the symmetry energy expansion and the peak value of the bulk viscosity

### v1

@@ V Conclusions
Parameterizing the EoS by expanding the symmetry energy around nuclear saturation density has become a standard technique in studying the EoS near saturation density. I have computed the susceptibilities of $npe$ matter - key components of the subthermal bulk viscosity - in terms of the symmetry energy expansion coefficients (Eqs 19a and 19b). I found that the susceptibility $A$, which describes the dependence of the pressure on the proton fraction, vanishes at saturation density if $J=L$, as this condition causes the proton fraction reaches a local maximum at saturation density. Furthermore, I derived an equation (Eq. 13) involving the symmetry energy and its derivative which can be solved to yield the densities where conformal points occur in the EoS.
@@
I apply the expression of the bulk viscosity in terms of the symmetry energy to the study of the r-mode instability window. The maximum stable spin rate of a neutron star as a function of temperature is expected to peak, and I find that for small values of $L$ and a relatively broad range of $K_{\text{sym}}$, this peak could lie below the Kepler frequency, meaning that a rapidly rotating neutron star containing only $npe$ matter could cool from extremely high temperatures and never pass through a temperature where the bulk viscosity would be high enough to stabilize the r-modes.
The entirety of this paper applies only to neutron stars constructed of $npe$ matter. A natural correction is to include muons. Muons add another term to the energy per baryon (Eq. 9), and their mass cannot be neglected. However, they are still a free Fermi gas, and don’t modify the symmetry energy expansion itself. More fundamentally, they add another equilibration channel, where not only does the proton fraction equilibrate by the Urca process, but the muon fraction equilibrates by the muon-Urca process (where the electron is replaced by a muon) and also by the decay of muons to electrons. This leads to, in principle, two peaks in the bulk viscosity instead of just one. However the peaks are not, in practice, separate and their interaction makes it difficult to have a simple expression for the maximum of the bulk viscosity [38, 39, 91, 49]. Recently, an effort was made to construct a symmetry energy expansion including strangeness [112, 113]. It would be interesting to explore whether conclusions about the hyperonic bulk viscosity [41, 114] can be made. Extensions of the work in this paper to suprathermal bulk viscosity would also be interesting, and of relevance to the saturation of the r-mode at large amplitudes [29]. Finally, ultimately the r-mode is a general-relativistic phenomenon, and yet all of the calculations of the mode amplitude profile I relied on were Newtonian. A relativistic study would be a natural upgrade [115].

### v3

@@ V Conclusions
Parameterizing the EoS by expanding the symmetry energy around nuclear saturation density has become a standard technique in studying the EoS near saturation density. I have computed the susceptibilities of npenpe matter - key components of the subthermal bulk viscosity - in terms of the symmetry energy expansion coefficients (Eqs 19a and 19b). I found that the susceptibility $A$, which describes the dependence of the pressure on the proton fraction, vanishes at saturation density if $J=L$, as this condition causes the proton fraction reaches a local maximum at saturation density. Furthermore, I derived an equation (Eq. 13) involving the symmetry energy and its derivative which can be solved to yield the densities where conformal points occur in the EoS.
@@
I apply the expression of the bulk viscosity in terms of the symmetry energy to the study of the r-mode instability window. The maximum r-mode-stable stable spin rate of a neutron star as a function of temperature is expected to peak, and I find that for small values of $L$ and a relatively broad range of $K_{\text{sym}}$, this peak could lie below the Kepler frequency, meaning that a rapidly rotating neutron star containing only npenpe matter could cool from extremely high temperatures and never pass through a temperature where the bulk viscosity would be high enough to stabilize the r-modes.
The entirety of this paper applies only to neutron stars constructed of npenpe matter. A natural correction is to include muons. Muons add another term to the energy per baryon (Eq. 9), and their mass cannot be neglected. However, they are still a free Fermi gas, and don’t modify the symmetry energy expansion itself. More fundamentally, they add another equilibration channel, where not only does the proton fraction equilibrate by the Urca process, but the muon fraction equilibrates by the muon-Urca process (where the electron is replaced by a muon) and also by the decay of muons to electrons. This leads to, in principle, two peaks in the bulk viscosity instead of just one. However the peaks are not, in practice, completely separate and their interaction makes it difficult to have a simple expression for the maximum of the bulk viscosity [38, 39, 96, 49]. Recently, an effort was made to construct a symmetry energy expansion including strangeness [117, 118]. It would be interesting to explore whether conclusions about the hyperonic bulk viscosity [41, 119] can be made. Extensions of the work in this paper to suprathermal bulk viscosity would also be interesting, and of relevance to the saturation of the r-mode at large amplitudes [29]. Finally, ultimately the r-mode is a general-relativistic phenomenon, and yet all of the calculations of the mode amplitude profile I relied on were Newtonian. A relativistic study would be a natural upgrade [120].